OpenStudy (anonymous):

Find the slope of the line tangent to the curve 2^x + 3x^2y+y = 1 at the point (0,1). Can anyone walk me through this?

OpenStudy (anonymous):

\[2^x+3x^2y+y = 1\] The slope of the line tangent to the curve of this?

zepdrix (zepdrix):

The line tangent to the curve is represented by the derivative function evaluated at the given point. So you'll need to take the derivative `implicitly`.

zepdrix (zepdrix):

So you have to apply a few different rules to take the derivative of each term. Remember how to deal with an exponential?\[\Large\rm \frac{d}{dx}2^x=?\]

OpenStudy (anonymous):

Not really, no.

zepdrix (zepdrix):

Mmm ok well maybe go review that at some point. You get the same thing back, but multiplied by the log of the base. Maybe you'll recall this derivative:\[\Large\rm \frac{d}{dx}e^x=e^x\]It's more accurately this (following the exponential rule for derivatives):\[\Large\rm \frac{d}{dx}e^x=e^x(\ln e)\]But since lne=1, we simply ignore it.

zepdrix (zepdrix):

For the second term, you need to apply the product rule, \[\Large\rm \left(3x^2y\right)'=\color{royalblue}{\left(3x^2\right)'}y+3x^2\color{royalblue}{\left(y\right)'}\]